Circular No. 6875

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998aq IN NGC 3982 G. M. Hurst, Basingstoke, England, reports the discovery by Mark

Armstrong, Rolvenden, of an apparent supernova (mag 14.9) on unfiltered CCD images obtained on Apr. 13.049, 13.080, and 13.109 UT with a 0.26m reflector in the course of the U.K. Nova/Supernova Patrol. The new star is located at α = 11^{h}56^{m}26^{s}.00, δ = +55^{o}07′38′′.8 (equinox 2000.0). R. Arbour, South Wonston, England, confirmed the presence of the object on unfiltered CCD exposures taken on Apr. 13.086 and 13.814 with a 0.30-m reflector, noting that SN 1998aq is 18′′ west and 7′′ north of the center of NGC 3982 (Hurst estimated the star’s magnitude as ∼ 14.5 on the latter exposure). Earlier patrol photos by Armstrong from Mar. 9 and Apr. 7 did not show the object; it is also absent from the Vickers CCD atlas, and there is no star to mag 19 in the USNO A1.0 catalogue.

CM DRACONIS H.-J. Deeg, Instituto de Astrofisica de Canarias; L. R. Doyle and J. M.

Jenkins, SETI Institute; and E. L. Martin, University of California, Berkeley, write: “The results of Guinan et al. on IAUC 6864 are inconsistent with our observational data. From 32 of our eclipse-minima timings obtained between May 1994 and Oct. 1996, we obtained residuals from the expected O − C values measured to better than 5.5 s; these O − C values were calculated from new epoch and period values for the CM Dra system (Deeg et al. 1998, submitted to A.Ap.). We find no evidence of significant periodic variations in the O−Cs, either at a period of 70.3 days or any other periods less than a limiting value of ∼ 3 yr (this maximum given by the length of our observational coverage); the standard deviation of our O − C values is 7.4 s, which makes any sinusoidal signal in the eclipse-minima times with amplitudes of > 10.5 s extremely unlikely (and amplitudes as high as 18 s certainly recognizable). We also report that the possible planetary transit eclipse of 0.08 mag reported on IAUC 6425 was an artifact produced by a dust speck in that night’s CCD frames; there are no indications in our database from 1994 to 1996 (617 hr) of any planetary transits with amplitudes larger than 0.01 mag. Our present phase coverage for the detection of coplanar planets around CM Dra ranges from 92 percent for 10-day periods to 64 percent for 30-day periods. From a work in progress (Doyle et al.), we can presently rule out coplanar planets in this data set around CM Dra larger than 3.0 R_{⊕} with orbital periods of less than ∼ 30 days at a confidence level of 90 percent.”

1998 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green